OpenStudy (anonymous):

Identify the maximum and minimum values of the function y = 8 cos x in the interval [-2π, 2π]. Use your understanding of transformations, not your graphing calculator.

OpenStudy (anonymous):

differentiate y'=-8sin(x) find critical values ( set the derivative equal to zero) -8sin(x)=0 sin(x)=0 x=0,pi, put in the function y=8cos(0)=8 y=8cos(pi)=-8 so maximium at x=0 is 8 minimium at x =pi is -8

OpenStudy (anonymous):

Where I am confused at is where did the y'=-8sin(x) come from?

OpenStudy (e.mccormick):

8 is a constant so it comes along for the ride. Then look at your rules for: \[\frac{ dy }{dx }\cos x\] Know that one by heart? If not, here are some proofs to study: http://www.math.com/tables/derivatives/more/trig.htm